OpenStudy (pottersheep):

(x+2) / (x-4) = (3x+1)/(x+2) How do I find x?

OpenStudy (lalaly):

first u do cross multiplication

OpenStudy (lalaly):

(x+2)(x+2)=(x-4)(3x+1)

OpenStudy (lalaly):

\[x^2+4x+4=3x^2-11x-4\]

OpenStudy (lalaly):

u know how to do the rest right?

OpenStudy (pottersheep):

Thanks ! Yes I do

OpenStudy (lalaly):

:D:D

OpenStudy (anonymous):

\[x=-(1/2), x=8\]

OpenStudy (anonymous):

(x+2) / (x-4) = (3x+1)/(x+2) find cross products (x+2)(x+2)=(x-4)(3x+1) \[x ^{2}+4x+4=3x ^{2}-11x-4\]\[2x ^{2}-15x-8=0\]\[(2x+1)(x-8)=0\]\[2x+1=0, x-8=0\]\[x=-1/2, x=8\] Source: http://www.mathskey.com